# Bidirectional Autoregressive Diffusion Model for Dance Generation

### v3

@@ 1 Introduction
Dance is a highly effective medium for expressing emotions, facilitating communication, and fostering social interaction. Nevertheless, the creation of new dances remains a formidable challenge, given the inherently expressive and freeform nature of dance movements. In order to generate new dance sequences that seamlessly integrate with both preceding motions and the musical context, various deep learning-based methods have been proposed [22, 29, 9, 31, 21, 37]. However, they typically treat dance generation as a matching problem. For example, some works [44, 53] rely on constructing a codebook for music-to-motion matching, which hinders their potential for creative dance generation. And they exclusively rely on past movements as the sole guide, neglecting the insights offered by future distributions. These methods also encounter limitations when attempting to generate dances from music based on user-defined constraints.
@@ 7 Conclusion
Supplementary Material
FID results.
As claimed in EDGE, the FID metrics present limitations since AIST++ test set does not thoroughly cover the train distribution. Even so, we also present the FID results in Table 4. We use the codes from Bailando. The FACT and Bailando results were reported from their original papers. Our method gets the best results except for the $\text{FID}_{\text{g}}$ of Bailando, demonstrating its capability to generate smoother sequences. Bailando focuses on organizing the movements, and $\text{FID}_{\text{g}}$ metric reflects the quality of choreography.
Method $\text{FID}_{\text{k}}\downarrow$ $\text{FID}_{\text{g}}\downarrow$
GT 17.10 10.60
FACT 35.31 22.11
Bailando 28.16 9.62
EDGE 31.23 24.71
Ours 27.02 20.24
Table 4: FID results of the different methods.
The reason for using $z_{k+1}$ in AR encoder.
According to the characteristics of the DDPM diffusion model, as the timestep t increases, the added noise becomes increasingly subtle. Consequently, the noised sequence bears a greater resemblance to the original one. $z_{k+1}$ progressively offers clearer insights about the future sequence. Also, from the FID results in Table 5, it can be seen that the model using a bidirectional encoder yields smoother results compared to the model using a unidirectional encoder.
Method $\text{FID}_{\text{k}}\downarrow$ $\text{FID}_{\text{g}}\downarrow$
EDGE 31.23 24.71
w/o Beat 27.60 20.94
Unidirection 28.24 22.28
w/o LID 28.31 22.51
Ours 27.02 20.24
Table 5: FID results of the different settings.
Details about diffusion process.
We employ DDPM to elaborate on the entire diffusion process, with a diffusion timestep set to 1000. Throughout the training, we randomly select the time step t in each iteration, effectively constraining the training time. For inference, we leverage DDIM for expediting sampling.
Inference spreed.
In our setting, EDGE requires 9.54 seconds to generate a 5-second dance sequence, while our method takes 10.45 seconds. Although our method incurs a slightly higher time cost, the difference is marginal.

### v4

@@ 1 Introduction
Dance is a highly effective medium for expressing emotions, facilitating communication, and fostering social interaction. Nevertheless, the creation of new dances remains a formidable challenge, given the inherently expressive and freeform nature of dance movements. In order to generate new dance sequences that seamlessly integrate with both preceding motions and the musical context, various deep learning-based methods have been proposed [22, 29, 9, 31, 21, 37]. However, they typically treat dance generation as a matching problem. For example, some works [56, 44, 53] rely on constructing a codebook for music-to-motion matching, which hinders their potential for creative dance generation. And they exclusively rely on past movements as the sole guide, neglecting the insights offered by future distributions. These methods also encounter limitations when attempting to generate dances from music based on user-defined constraints.